The maen travel time of light Puerto princesa city -iloilo is said to be 1.5hours .A certain airfare company recorded 20 different flights actual travel time.the mean is 2 hours with a standard deviation of 0.6hours .







1
Expert's answer
2022-06-06T23:48:50-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=1.5"

"H_1:\mu\not=1.5"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=19" and the critical value for a two-tailed test is "t_c =2.093024."

The rejection region for this two-tailed test is "R = \{t:|t|>2.093024\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{2-1.5}{0.6/\sqrt{20}}=3.7268"


Since it is observed that "t=3.7268>2.093024=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=19" degrees of freedom, "t=3.7268" is "p= 0.001426," and since "p= 0.001426<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 1.5, at the "\alpha = 0.05" significance level.